# ForeSight: A Predictive-Scheduling Deterministic Database

### v1

@@ Abstract.
[…]
@@ 1. Introduction
(3) Ineffective reordering for transaction scheduling Existing deterministic databases schedule transactions inefficiently due to ineffective reordering strategies. Optimistic deterministic databases (Lu et al., 2020) attempt to reorder transactions using coarse rules (e.g., aborting transactions that depend on earlier ones) without analyzing whether dependencies actually form cycles, leading to unnecessary aborts. Pessimistic deterministic databases (Faleiro and Abadi, 2015; Faleiro et al., 2017) build fine-grained dependency graphs, but their graph-based reordering is computationally expensive, making them impractical at large batch sizes or high throughput.
@@
Together, these three techniques instantiate predictive scheduling, addressing the key challenges of existing deterministic databases.
@@
  • Extensive evaluations demonstrate that ForeSight consistently outperforms state-of-the-art deterministic databases. ForeSight achieves up to $2\times$ throughput improvement on skewed YCSB workloads and maintains superior stability under high-conflict TPC-C scenarios (Section 7).
@@ 4.2. The ASPN Model
The Background of SPN. Sum-Product Network (SPN) is a probabilistic model that factorizes a joint distribution using sum and product nodes (Hilprecht et al., [n.d.]), enabling linear-time inference by partitioning attributes under independence assumptions and decomposing heterogeneous data regions. The main advantage of SPN is that it allows efficient probability computation—complex queries can be answered in time that scales with the network size, making it practical for large-scale data analysis. However, real-world database workloads often violate independence assumptions, and excessive partitioning can result in large, inefficient structures. To address these limitations, the factorize-split-sum-product network (FSPN) (Zhu et al., 2021) extends SPN to accommodate the demands of database workloads. However, its design remains focused on estimating query result sizes rather than supporting the efficient predicate intersection required for conflict prediction.
To support accurate and efficient conflict prediction, we propose Association Sum-Product Network (ASPN), a tree-structured probabilistic model that extends SPN by applying conditional decomposition to groups of highly correlated attributes. By building on SPN, ASPN keeps the benefit of efficient probability computation while adding the ability to capture correlations that SPN cannot handle. Unlike SPN, which relies on independence assumptions and suffers from expressiveness limitations, ASPN selectively models joint distributions for correlated subsets, enabling compact yet expressive representations. This design is particularly effective for conflict prediction, where the goal is to determine whether the conjunction of two transaction predicates intersects a non-empty region of the data space. Moreover, ASPN avoids the common overestimation of overlaps caused by assuming attribute independence, which often leads to false positives and unnecessary aborts. By learning attribute correlations from data, ASPN estimates predicate intersections more precisely and filters out overlaps corresponding to empty or low-probability regions. The improved accuracy in conflict prediction enables finer-grained scheduling, reducing false detections and wasted work, and thereby improving throughput under high contention.
@@ 5. Multi-Version-Based Optimization
Deterministic systems often suffer from frequent aborts under high contention due to aggressive validation rules that reject transactions with potential conflicts, leading to reduced throughput and wasted computation. To address this, ForeSight introduces a multi-version-based optimization that relaxes commit constraints while preserving correctness, enabling more transactions to commit successfully. This section presents the design of our multi-version-based execution process (Section 5.1), the fallback strategy for transactions that fail initial validation (Section 5.2), and a formal proof of determinism and serializability (Section 5.3).
@@ 5.1. Multi-Version-Based Execution Process
In ForeSight, all transactions are executed on a consistent snapshot, avoiding synchronization during execution. Instead of modifying records in place, transactions buffer their updates as new versions, which remain invisible until commit. This deferred write strategy eliminates mid-execution conflicts and simplifies validation. The multi-version storage organizes each record as a version chain, with the head pointing to the latest visible version. Each version node stores the value, a pointer to its predecessor, and metadata such as a commit timestamp or transaction ID (TID). At the start of a batch, the head points to a base version representing the snapshot. During the commit phase, transactions append new version nodes for their writes, set timestamps to their TIDs, and atomically update the chain head. This organization ensures that writes produce new versions rather than modifying records in place. Reads always access the snapshot-visible versions, enabling conflict-free execution in the execution phase.
@@ 5.2. Fallback Strategy
To address these limitations, ForeSight adopts a multi-version-based fallback strategy. Fallback transactions run on the version chains generated at the end of the execution phase, which already includes updates from all executed transactions. Instead of modifying records in place, each transaction appends its writes as new versions in the version chains, tagged with its TID and kept invisible until commit. When reading, a transaction retrieves the latest version whose timestamp precedes its own TID. Transactions in the fallback phase can execute in parallel, while their commits are finalized in deterministic timestamp order so that each incorporates all prior commits, preventing inconsistent reads. Following BOHM’s “reads don’t block writes” principle (Faleiro and Abadi, 2015; Tu et al., 2013), this design eliminates global lock acquisition, avoids blocking among threads, and removes coordination overhead while preserving serializability.
Figure 3 overviews the multi-version-based optimization in ForeSight. A batch of transactions executes with snapshot reads guided by conflict-aware planning, while unsafe ones enter a fallback phase that pre-inserts writes, reads from version chains, and validates before commit. Figures 3(a)–(b) give an example: transactions $T_{3}$ and $T_{6}$ abort due to RAW/WAW conflicts on $T_{1}$ and $T_{2}$, while $T_{7}$, $T_{8}$, and $T_{10}$ abort on other keys. With reordering (Section 6), $T_{4}$ can still commit despite a RAW dependency. In fallback execution, write transactions extend version chains sequentially, while reads traverse chains for visible versions. Multi-versioning ensures that even transactions with conflicts can proceed without blocking, with validation deferred to commit.
@@ 5.3. Determinism and Serializability
Determinism: Given a batch of transactions and a database snapshot, ForeSight assigns fixed transaction IDs and executes all transactions on the same snapshot. Each transaction’s commit decision is based solely on its read/write sets and the presence of RAW dependencies with lower-TID transactions. Because the input, snapshot, and validation rule (Rule 1) are identical across all replicas, the commit outcome of each transaction is deterministic.
Serializability: The following proof is by contradiction. Assume there exists a committed transaction sequence $\cdots\rightarrow T_{i}\rightarrow\cdots\rightarrow T_{j}\rightarrow\cdots$ and the result produced by ForeSight is not serializable. Since transactions are executed concurrently and then validated for commit, two possible cases may cause non-serializability: (1) Transaction $T_{j}$ read a version written by $T_{i}$ (RAW), and (2) Transaction $T_{j}$’s updates were overwritten by $T_{i}$’s updates (WAW). For (1), all transactions in a batch execute on the same consistent snapshot, and reads only access committed versions from prior rounds. Thus, $T_{j}$ cannot read a version written by $T_{i}$ within the same batch. If it did, it would violate snapshot isolation—contradicting ForeSight’s execution semantics. For (2), ForeSight employs multi-versioning, so each transaction appends a new version rather than overwriting existing data. Therefore, $T_{i}$ cannot overwrite $T_{j}$’s updates, and WAW does not cause write loss or inconsistency. By Rule 1, transactions with RAW dependencies on prior transactions are rejected, and WAW dependencies are safely handled via version chains. Both cases lead to a contradiction. Hence, the execution must be conflict-serializable. ∎
@@ 6.1. A Motivating Example
Figure 4. Comparison of reordering strategies and the validation rules of ForeSight.
The ordering of input transactions is the key to scheduling in deterministic databases, as it determines the commit order and affects the overall throughput (Lu et al., 2020). Figure 4 illustrates two representative reordering strategies.
@@
(2) With Dependency Graph: The FVS-Greedy Algorithm (FGA) (Ding et al., 2018) constructs the dependency graph after executing the batch, detects strongly connected components (SCCs) (Nuutila and Soisalon-Soininen, 1994), and breaks cycles by aborting the transaction with the highest weight (based on in/out degrees) in each multi-node SCC. Here, a feedback vertex set (FVS) refers to the minimal set of vertices whose removal eliminates all cycles.
We presented a design where any transaction satisfying Rule 1 can commit (Section 5). According to the reordering principle of Aria (Lu et al., 2020), the abort rate can be reduced by transforming RAW dependencies into WAR dependencies. With snapshot reads, a RAW dependency always causes an abort because the read cannot observe a future write, while reordering it into a WAR dependency still yields serializable where the read completes before the write. Applying such reordering enables Rule 1 to be relaxed to Rule 2.
@@ 6.2. Algorithm Design
Given $M$ and the $RAW$ dependency dictionary, the algorithm outputs an abort set $S\subseteq{T_{1},\dots,T_{N}}$. The method proceeds by identifying all dependency paths that involve RAW edges and selecting the transactions that occur most frequently on these paths.
@@ 6.3. Correctness and Effectiveness Analysis
$\Pr[T_{i}]$ $=\left[1-(1-\Pr[\neg\text{RAW}_{i}])(1-\Pr[\neg\text{WAR}_{i}])\right]$
$\quad\cdot\left[1-(1-\Pr[\neg\text{RAW}_{i}])(1-\Pr[\neg\text{WAW}_{i}])\right].$
@@ 7.1. Experimental Setup
(c) Throughput scalability on TPC-C with varying numbers of partitions
@@
Workload Configurations. To evaluate the dependency prediction accuracy, we generate 1,000 single-table queries on the GAS dataset and both 1,000 single-table and 1,000 multi-table queries on TPCH, IMDB, and TPC-C. All queries are randomly generated and contain at least one range predicate.
The protocol performance is assessed on TPC-C and YCSB, which allow efficient pre-computation of read/write sets and are widely used in deterministic database evaluations. In TPC-C, transactions operate on ten warehouse partitions with ten operations per transaction, and the number of partitions matches the number of worker threads. YCSB includes two variants, $A$ with a balanced 50% read and 50% update mix, and $B$ with a read-heavy 95% read and 5% update mix, using a Zipfian distribution to simulate skew. Performance results are based on YCSB-A, except PWV, where PWV-A/B denote runs on YCSB-A/B due to early write visibility effects. Each partition stores 40,000 keys. Batch sizes are 1,000 for YCSB and 500 for TPC-C.
@@ 7.2. Evaluation of Protocol Performance
In this section, we compare the performance of ForeSight against state-of-the-art deterministic databases and a traditional primary-backup system PB using the YCSB and TPC-C benchmarks.
@@ 7.3. Effectiveness of Conflict Prediction
We evaluate the conflict prediction algorithm on TPCH and TPC-C workloads, covering both single-table and multi-table queries. The ASPN model is built on 10,000 sampled records and tested on 200 queries. For comparison, we adapt the Supervised Machine Learning method as baseline (Sheng et al., 2019; Nasteski, 2017), which hashes transaction features and applies logistic regression to predict conflicts (LaValley, 2008). The model is trained on 800 labeled queries and tested on a separate set of 200 queries. Both models predict conflicts in batch mode. Evaluation metrics include accuracy (acc), precision (prec), recall, prediction time ($T_{p}$), and execution time ($T_{e}$), where $T_{p}$ measures the time required for conflict prediction, and $T_{e}$ measures the time required to obtain conflicts through actual query execution.
@@
As shown in Table 1, we observe that ASPN consistently outperforms baseline across all settings. For single-table queries, ASPN achieves near-perfect accuracy (99.9% on TPCH and 95.6% on TPC-C) and precision (100% and 92.7%, respectively), while maintaining high recall. In contrast, baseline exhibits significantly lower accuracy (72.3%) and precision (73.5%) on TPCH and fails to maintain high recall on multi-table queries. This indicates its limited capability to capture complex transactional patterns across joins or conditional logic. For multi-table queries, ASPN retains strong performance (91.3% and 95.2% accuracy; 88.4% and 93.9% precision on TPCH and TPC-C, respectively), while baseline degrades to near-random precision and recall, demonstrating its inability to generalize beyond simple access patterns. In terms of prediction time, ASPN achieves effective conflict prediction, taking 0.437 s and 1.123 s on single-table TPCH and TPC-C, and 0.400 s and 1.415 s on multi-table queries. Compared with query execution, which requires 19.7 s and 97.1 s for TPCH and 51.1 s and 96.7 s for TPC-C, the prediction overhead is negligible, accounting for only 2.2% and 0.4% of TPCH execution time and remaining below 3% on TPC-C. This demonstrates that ASPN achieves high prediction accuracy while maintaining very low inference cost relative to query execution.
@@
We next evaluate the robustness of the ASPN model under different training sample sizes, ranging from 10,000 to 50,000 in increments of 10,000, across both single-table and multi-table workloads. Figure 9 shows that in single-table scenarios, precision remains consistently high as the sample size grows. Most workloads maintain over 94% precision, with IMDB being the exception at around 75% due to more complex query patterns. While in multi-table scenarios, precision is similarly stable for most workloads. TPCH shows steady improvement with larger training sets, reflecting the benefits of more samples in capturing complex inter-table relationships. Training time increases with sample size in all cases and correlates with workload complexity: GAS has the shortest training time (14.62 s–40.58 s) due to simpler correlations, while multi-table TPCH is the longest (63.05 s–271.38 s). Figure 10 shows that prediction time also grows roughly linearly with sample size but remains under 2 seconds for all workloads. This low overhead, combined with stable or improving precision, demonstrates that ASPN scales well to large datasets and remains suitable for real-time conflict prediction in high-throughput systems.

### v2

@@ Abstract.
Deterministic databases enable scalable replicated systems by executing transactions in a predetermined order. However, existing designs fail to capture transaction dependencies, leading to insufficient scheduling, high abort rates, and poor resource utilization. By addressing these challenges with lightweight conflict prediction and informed scheduling, we present ForeSight, a high-performance deterministic database system. Our system has three core improvements: (1) We design an Association Sum-Product Network to predict potential transaction conflicts, providing the input for dependency analysis without pre-obtained read/write sets. (2) We enhance the storage engine to integrate multi-version-based optimization, improving the execution process and fallback strategy to boost commit rates and concurrency. (3) We propose a matrix two-pass forward scan algorithm that performs dependency analysis to generate conflict-aware schedules, significantly reducing scheduling overhead. Experimental results on multiple benchmarks show that ForeSight achieves up to 2 times higher throughput on skewed workloads and maintains strong performance under contention, demonstrating that predictive scheduling substantially improves deterministic database scalability.
@@ 1. Introduction
(3) Ineffective reordering for transaction scheduling. Existing deterministic databases schedule transactions inefficiently due to ineffective reordering strategies. Optimistic deterministic databases (Lu et al., 2020) attempt to reorder transactions using coarse rules (e.g., aborting transactions that depend on earlier ones) without analyzing whether dependencies actually form cycles, leading to unnecessary aborts. Pessimistic deterministic databases (Faleiro and Abadi, 2015; Faleiro et al., 2017) build fine-grained dependency graphs, but their graph-based reordering is computationally expensive, making them impractical at large batch sizes or high throughput.
@@
Together, these three techniques enable predictive scheduling, addressing the key challenges of existing deterministic databases.
@@
  • Extensive evaluations demonstrate that ForeSight consistently outperforms state-of-the-art deterministic databases. ForeSight achieves up to $2\times$ throughput improvement on skewed YCSB workloads and maintains superior stability under high-conflict TPC-C workloads (Section 7).
@@ 4.2. The ASPN Model
We build on the Sum-Product Network (SPN) as the foundation for ASPN. An SPN is a probabilistic model that factorizes a joint distribution using sum and product nodes (Hilprecht et al., [n.d.]), enabling linear-time inference by partitioning attributes under independence assumptions and decomposing heterogeneous data regions. The main advantage of SPN is its efficient and deterministic probability computation, since complex queries can be answered in time that scales with the network size, and once the structure and parameters are fixed, identical inputs always produce identical outputs. Such stability makes SPN appropriate as a deterministic inference component, where consistent results are essential to ensure determinism. However, SPN is not well-suited for database workloads, as real queries rarely satisfy the strong independence assumptions it relies on, and excessive partitioning often produces large structures. The factorize-split-sum-product network (FSPN) (Zhu et al., 2021) partially addresses these issues, but its design is primarily aimed at cardinality estimation rather than the predicate intersection required for conflict prediction.
To overcome these limitations, we propose the Association Sum-Product Network (ASPN), a tree-structured probabilistic model that applies conditional decomposition to groups of highly correlated attributes. This design preserves SPN’s efficiency and determinism while selectively modeling joint distributions for correlated subsets, yielding compact yet expressive representations. ASPN is particularly effective for conflict prediction, where the task is to determine whether two transaction predicates intersect a non-empty region of the data space. By capturing attribute correlations, it avoids the overestimation caused by independence assumptions, thereby reducing false positives and unnecessary aborts. This enables predictive scheduling that reduces transaction aborts and resource waste, ultimately improving throughput under high contention. Furthermore, since ASPN construction depends only on the given training samples and produces a fixed model once built, its inference yields consistent results. Given the same data, both the model and its predictions remain unchanged, making ASPN especially suitable for conflict prediction in deterministic databases.
@@ 5. Multi-Version-Based Optimization
Deterministic systems often suffer from frequent aborts under high contention due to aggressive validation rules that reject transactions with potential conflicts, leading to reduced throughput and wasted computation. To address this, ForeSight introduces a multi-version-based optimization that relaxes commit constraints while preserving correctness, enabling more transactions to execute successfully. This section presents the design of our multi-version-based execution process (Section 5.1), the fallback strategy for transactions that fail initial validation (Section 5.2), and a formal proof of determinism and serializability (Section 5.3).
@@ 5.1. Multi-Version-Based Execution Process
In ForeSight, all transactions are executed on a consistent snapshot, avoiding synchronization during execution. Instead of modifying records in place, transactions buffer their updates as new versions, which remain invisible until commit. This deferred write strategy eliminates mid-execution conflicts and simplifies validation. The multi-version storage organizes each record as a version chain, with the head pointing to the latest visible version. Each version node stores the value, a pointer to its predecessor, and metadata such as a commit timestamp or transaction ID (TID). At the start of a batch, the head points to a base version representing the snapshot. During the commit phase, transactions append new version nodes for their writes, set timestamps to their TIDs, and atomically update the chain head. This organization ensures that write operations produce new versions rather than modifying records in place. Reads always access the snapshot-visible versions, enabling conflict-free execution in the execution phase.
@@ 5.2. Fallback Strategy
To address these limitations, ForeSight adopts a multi-version-based fallback strategy. Fallback transactions run on the version chains generated at the end of the execution phase, which already include updates from all executed transactions. Instead of modifying records in place, each transaction appends its writes as new versions in the version chains, tagged with its TID and kept invisible until commit. When reading, a transaction retrieves the latest version whose timestamp precedes its own TID. Transactions in the fallback phase can execute in parallel, while their commits are finalized in deterministic timestamp order so that each incorporates all prior commits, preventing inconsistent reads. Following BOHM’s “reads don’t block writes” principle (Faleiro and Abadi, 2015; Tu et al., 2013), this design eliminates global lock acquisition, avoids blocking among threads, and removes coordination overhead while preserving serializability.
Figure 3 overviews the multi-version-based optimization in ForeSight. A batch of transactions executes with snapshot reads guided by conflict-aware planning, while unsafe ones enter a fallback phase that pre-inserts writes, reads from version chains, and validates before commit. Figures 3(a)–(b) give an example: transactions $T_{3}$ and $T_{6}$ abort due to RAW/WAW conflicts on $T_{1}$ and $T_{2}$, while $T_{7}$ aborts on other keys, and $T_{8}$ and $T_{10}$ are early deferred by ASPN-based predictive scheduling. With the reordering strategy introduced in Section 6, the RAW dependency of $T_{4}$ is reordered into a WAR dependency, so $T_{4}$ can still commit. In the fallback phase, write transactions extend version chains sequentially, while reads traverse chains for visible versions. Multi-versioning ensures that even transactions with conflicts can proceed without blocking, with validation deferred to commit.
@@ 5.3. Determinism and Serializability
Determinism: Given a batch of transactions and a database snapshot, ForeSight assigns fixed transaction IDs and then uses ASPN to filter them, executing the filtered transactions on the same snapshot. Each transaction’s commit decision is based solely on its read/write sets and the presence of RAW dependencies with lower-TID transactions. Since ASPN is deterministic (Section 4), the same batch of transactions always produces the same filtering result, keeping the input consistent across replicas. Combined with the identical snapshot and validation rule (Rule 1), this guarantees that all replicas maintain deterministic execution.
Serializability: The following proof is by contradiction. Assume there exists a committed transaction sequence $\cdots\rightarrow T_{i}\rightarrow\cdots\rightarrow T_{j}\rightarrow\cdots$ and the result produced by ForeSight is not serializable. With concurrent execution and commit validation, two cases may cause non-serializability: (1) Transaction $T_{j}$ read a version written by $T_{i}$ (RAW), and (2) Transaction $T_{j}$’s updates were overwritten by $T_{i}$’s updates (WAW). For (1), all transactions in a batch execute on the same consistent snapshot, and reads only access committed versions from prior rounds. Thus, $T_{j}$ cannot read a version written by $T_{i}$ within the same batch. If it did, it would violate snapshot isolation, contradicting ForeSight’s execution semantics. For (2), ForeSight employs multi-versioning, so each transaction appends a new version rather than overwriting existing data. Therefore, $T_{i}$ cannot overwrite $T_{j}$’s updates, and WAW does not cause write loss or inconsistency. By Rule 1, transactions with RAW dependencies on prior transactions are rejected, and WAW dependencies are safely handled via version chains. Both cases lead to a contradiction. Hence, the execution must be conflict-serializable. ∎
@@ 6.1. A Motivating Example
Figure 4. Illustration of two reordering strategies and the validation rules of ForeSight
The ordering of input transactions is the key to scheduling in deterministic databases, as it determines the commit order and affects the overall throughput (Lu et al., 2020). Figure 4 illustrates two representative reordering strategies and the validation rules of ForeSight.
@@
(2) With Dependency Graph: The FVS-Greedy Algorithm (FGA) (Ding et al., 2018) constructs the dependency graph after executing the batch, detects strongly connected components (SCCs) (Nuutila and Soisalon-Soininen, 1994), and breaks cycles by aborting the transaction with the highest weight (based on in/out degrees) in each multi-node SCC. This approach is based on the notion of a feedback vertex set (FVS), which refers to the minimal set of vertices whose removal eliminates all cycles.
We presented a design where any transaction satisfying Rule 1 can commit (Section 5). According to the reordering principle of Aria (Lu et al., 2020), the abort rate can be reduced by transforming RAW dependencies into WAR dependencies. Under snapshot reads, a RAW dependency leads to an abort because the read cannot see future writes. If reordered into a WAR dependency, the execution remains serializable, with the read finishing before the write. Applying such reordering enables Rule 1 to be relaxed to Rule 2.
@@ 6.2. Algorithm Design
Given $M$ and the dependency dictionary $RAW$, the algorithm outputs an abort set $S\subseteq{T_{1},\dots,T_{N}}$. The method proceeds by identifying all dependency paths that involve RAW edges and selecting the transactions that occur most frequently on these paths.
@@ 6.3. Correctness and Effectiveness Analysis
$\Pr[T_{i}]=1-(1-\Pr[\neg\text{RAW}_{i}])\cdot(1-\Pr[\neg\text{WAR}_{i}]\cdot\Pr[\neg\text{WAW}_{i}])$
@@ 7.1. Experimental Setup
(c) Throughput under TPC-C with varying numbers of partitions
@@
Workload Configurations. We evaluate protocol performance and the effectiveness of reordering algorithms on TPC-C and YCSB, which support efficient pre-computation of read/write sets and are widely used in deterministic database studies. In TPC-C, transactions operate on ten warehouse partitions with ten operations per transaction, and the number of partitions equals the number of worker threads. YCSB includes two variants, $A$ with a balanced 50% read/50% update mix and $B$ with a read-heavy 95% read/5% update mix, both using a Zipfian distribution to simulate skew. Performance results are based on YCSB-A, except for PWV, where PWV-A/B denote runs on YCSB-A/B due to early write visibility effects. Each partition stores 40,000 keys, with batch sizes of 1,000 for YCSB and 500 for TPC-C. We also evaluate conflict prediction using 1,000 single-table queries on GAS and both 1,000 single-table and 1,000 multi-table queries on TPCH, IMDB, and TPC-C; all queries are randomly generated and contain at least one range predicate.
@@ 7.2. Evaluation of Protocol Performance
In this section, we compare the performance of ForeSight against state-of-the-art deterministic databases and the traditional primary-backup system PB, using the YCSB and TPC-C benchmarks.
@@ 7.3. Effectiveness of Conflict Prediction
We evaluate the conflict prediction algorithm on TPCH and TPC-C workloads, covering both single-table and multi-table queries. The ASPN model is built on 10,000 sampled records and tested on 200 queries. As a baseline, we employ the learned abort prediction method proposed in (Sheng et al., 2019). This method applies supervised learning by hashing transaction features and using logistic regression to predict conflicts (Nasteski, 2017; LaValley, 2008). The model is trained on 800 labeled queries and tested on a separate set of 200 queries. Both models predict conflicts in batch mode. Evaluation metrics include accuracy (acc), precision (prec), recall, prediction time ($T_{p}$), and execution time ($T_{e}$), where $T_{p}$ measures the time required for conflict prediction, and $T_{e}$ measures the time required to obtain conflicts through actual query execution.
@@
As shown in Table 1, we observe that ASPN consistently outperforms baseline across all settings. For single-table queries, ASPN achieves near-perfect accuracy (99.9% on TPCH and 95.6% on TPC-C) and precision (100% and 92.7%, respectively), while maintaining high recall. In contrast, baseline exhibits significantly lower accuracy (72.3%) and precision (73.5%) on TPCH and fails to maintain high recall on multi-table queries. This indicates its limited capability to capture complex transactional patterns across joins or conditional logic. For multi-table queries, ASPN retains strong performance (91.3% and 95.2% accuracy; 88.4% and 93.9% precision on TPCH and TPC-C, respectively), while baseline degrades to near-random precision and recall, demonstrating its inability to generalize beyond simple access patterns. In terms of prediction time, ASPN achieves effective conflict prediction, taking 0.437 s and 1.123 s on single-table TPCH and TPC-C, and 0.400 s and 1.415 s on multi-table queries. Compared with query execution, which requires 19.7 s and 97.1 s for TPCH and 51.1 s and 96.7 s for TPC-C, the prediction overhead is negligible, accounting for only 2.2% and 0.4% of TPCH execution time and remaining below 3% on TPC-C. This demonstrates that ASPN achieves high prediction accuracy while maintaining very low inference cost compared to query execution.
@@
We next evaluate ASPN with training sample sizes ranging from 10,000 to 50,000 across both single-table and multi-table workloads. As shown in Figure 9, training costs increase with sample size and correlate with workload complexity, ranging from 14.6 to 40.6 seconds for GAS to 28.6 to 106.6 seconds for TPCH in single-table workloads. Precision remains largely stable ($\geq$94%) across workloads, except for IMDB, which is lower ($\sim$75%) due to its complex query patterns. Although prediction accuracy may vary across workloads, ASPN inference remains deterministic with identical samples producing identical outputs. Combined with the reorder-validation safeguard, this property guarantees execution consistency and determinism. Figure 10 shows that prediction time grows roughly linearly with sample size but remains under 2 seconds for all workloads. This low overhead, combined with stable or improving precision, demonstrates that ASPN scales well to large datasets and remains suitable for real-time conflict prediction in high-throughput systems.